an electrical company claims that an average life of bulbs it manufactures is 1200 hours with standard deviation of 250 hours.If a random sample of 100 bulbs is chosen, what is the probability of the sample mean will be:




a.Greater than 1150 hours?



b.Less than 1250 hours?



c.Between 1150 and 1250 hours?​

1
Expert's answer
2022-04-18T17:00:37-0400

"\mu=1200 \\

\sigma = 250 \\

n=100"

a.

"P(\bar{x} > 1150) = 1 -P(\bar{x} < 1150) \\

= 1 -P(Z< \frac{1150-1200}{250 / \sqrt{100}} )\\

= 1 -P(Z< -2) \\

= 1 -0.0227 \\

= 0.9773"

b.

"P(\bar{x} < 1250) = P(Z < \frac{1250-1200}{250 / \sqrt{100}}) \\

= P(Z< 2) \\

= 0.9772"

c.

"P(1150 < \bar{x} < 1250) = P(\bar{x} < 1250) -P(\bar{x} < 1150) \\

= P(Z < \frac{1250-1200}{250 / \sqrt{100}}) -P(Z < \frac{1150-1200}{250 / \sqrt{100}}) \\

= P(Z< 2) -P(Z< -2) \\

= 0.9772 -0.0227 \\

= 0.9545"